# From tests to effect sizes: Quantifying uncertainty and statistical variability in multilingual and multitask NLP evaluation benchmarks

### v1

@@ 4.2 Resampling for model comparison
We seek to quantify the uncertainty in the XQUAD F1 scores in each language as well as aggregated across languages. Specifically, we compute pairwise model differences on each language as well as between aggregate scores computed using the arithmetic mean, geometric mean and median. We also compute rankings as point estimates. Even for the arithmetic mean, it is not possible to do it in closed form since we have no estimate of the covariance between the models. Instead, we estimate the difference of the models using the following parametric bootstrap resampling procedure. Given a language, we first take an estimate of its each model’s average performance on it by averaging the $S=5$ observed values. We then use estimate of the within-language SD, $\eta_{ml}$, to get a randomly resampled performance score for each model. We estimate $\eta_{ml}$ by first computing the standard deviation of scores across seeds and then use the average bootstrap SD $\tau_{ml}$ across the $S=5$ values computed by lm-eval-harness to form an aggregate SD estimate of $\eta_{ml}$. We then sample a noise term using a random draw from $\varepsilon\sim\mathcal{N}(0,\eta_{ml}^{2})$ yielding a resampled score $x+\varepsilon$ with variance $\eta_{ml}^{2}=\sigma_{ml}^{2}+\tau_{ml}^{2}$. While the score itself may not be plausibly Gaussian, we feel that modeling the noise terms as such is reasonable, as we observe mostly small score deviations across both seeds and bootstrap resamples. We repeat this process $R=1000$ times. For each replication, we compute a performance difference on each language, and an aggregated score of performance scores across languages, using the arithmetic mean, geometric mean and median. We also rank the models based on the randomly replicated performance scores. The results of this procedure are summarized in several tables below. Table 2 summarizes pairwise differences between models and their associated SDs. Table 3 gives the rank distribution of each model using each of the summary statistics. While TowerInstruct and aya reliably place first and last regardless of metric, the rank distributions of Clarus-7b and gemma2-9b can vary significantly depending on which aggregation function is used.
@@
5 Task 2: LLM-based machine translation
@@ 5.1 Variance components
The lm-eval-harness library we use to run our experiments automatically computes bootstrapped standard errors for BLEU, ChrF++ and COMET. In addition to the bootstrap SE, we incorporate model-side uncertainty by sampling 5 translation hypotheses for each source sentence from each model. The summarized results are shown in Table 8 in the appendix. A detailed variance decomposition of the within-language variation is shown in Table 11 in the appendix. As with XQUAD, the between-language variability ($\nu$) is much larger than either the model-side or bootstrap uncertainty. Most variance components are also less than 1.0 BLEU or ChrF++ points.
@@ 5.2 Resampling for model comparison
Like with XQUAD, we also wish to quantify the uncertainty in the performance differences between the two models on each language. Given observed BLEU/ChrF++/COMET scores for each language we can easily compute a point estimate by subtraction. We quantify the uncertainty in this point estimate by resampling as with XQUAD. The estimated differences are displayed in Table 7. The most obvious observation is that Clarus-7B clearly underperforms relative to aya-expanse-8b and TowerInstruct-7B. The latter two models perform more similarly to each other and significant differences were found in 6 out of 16 translation using BLEU. Out of the 6 significant differences, TowerInstruct-Mistral-7B beat aya-expanse-8b 4 out of 6 times. When using ChrF++, significant differences were found in 10 out of 16 translation tasks, with TowerInstruct-Mistral-7B as the winner in all of them.
@@ 6 Task 3: NER
As a third task, we show how resampling can be used for model comparison in multilingual NER when finetuning pretrained language models. Our analysis is based on the OpenNER 1.0 multilingual NER benchmark (Palen-Michel et al., 2024) which consists of 61 unique datasets covering a total of 51 languages . Since we reanalyze the results of the original paper, we use the same models: mBERT and XLM-R. For each model architecture, we vary the training data according to two conditions on each language-dataset pair: (i) individual, where each model is finetuned with data from only a given training set and (ii) concatenated, where each model is finetuned with a concatenated and downsampled version of all the training data included in OpenNER.
In addition to the original results, we construct $B=100$ bootstrap replicated data sets and reuse the same predictions from the original test set. This gives us a total of $R=SB=1000$ replicated scores for each of the 61 datasets.
@@ 6.1 Variance components via resampling
Like with QA and MT, we investigate how resampling-based inference can be used to better make sense of experimental variation by estimating the variance components due to model- and data-side uncertainty for each of the 61 languages pairs that comprise the overall benchmark. A concise display of inferences for $\nu$, $\sigma$ and $\tau$ are shown in Table 4. A more detailed table of estimates is shown in the Appendix in Table 12.
@@
Glot500 $1.59_{(2.51)}$ $0.61_{(0.39)}$
$\sigma_{l}$ mBERT $1.15_{(2.55)}$ $0.81_{(0.45)}$
XLM-R $1.24_{(3.53)}$ $0.64_{(0.42)}$
Glot500 $27_{(0.95)} 1.1$$17_{(0.79)}$
$\tau_{l}$ mBERT $1.36_{(0.86)}$ $1.47_{(0.91)}$
XLM-R $23_{(0.84)} 1.1$$18_{(0.79)}$
Table 4: Summary of variance components for NER. The three groups of rows, $\nu$, $\sigma$ and $\tau$ correspond to between-language, seed and bootstrap SDs. Subscripts indicate standard deviations across languages.
A few general observations are apparent. First, both seed and bootstrap variability tend to be less than 1.0 F1 points in most cases. Second, the average total within-language variability ($\eta_{l}$) is approximately 1.5 to 2 F1 points, suggesting that any deviations less than 3-4 F1 points will be statistically insignificant on any individual dataset.
When we compare models in the “individual” versus the “concatenated” conditions, there is a distinct decrease in the variability due to random seeds in the “concatenated” condition. This suggests that smaller languages may indeed benefit from larger finetuning datasets. At the same time, the variability due to bootstrapped test sets remains relatively constant, which suggests that larger finetuning sets do not help much in terms of dataset-to-dataset generalization.
@@ 6.2 Resampling for model comparison
To approximate quantities other than variance components, we resample the $R=SB=1000$ replications within each language. Unlike our QA and MT experiments, we do this entirely nonparametrically. We also show how the between-language score variability ($\nu$) can be incorporated into the analysis by also resampling what languages are considered when estimating overall performance. Our analysis based on effect sizes in Table 5 shows that underestimating variation may lead to an inflated sense of statistical significant differences between models results than when using randomization.
Average/overall performance
[…]
To estimate the standard error of each estimator, we sample $R=5000$ draws from the approximate joint distribution where we draw independently from the empirical distribution of scores for each language $l$. We conduct two separate simulations: one where the $L=61$ languages are kept constant and another one where they are randomly resampled with replacement. To compute an interval we use the Gaussian-inspired [^2] This is based on the property that approximately 95% of the variation lies within within $M=1.96\approx 2$ standard deviations of the point estimate. interval of $\text{estimate}\pm 2\text{se}$ Alternatively, we also use a percentile interval 2.5% and 97.5% percentiles of each estimator, computed over resampled draws. A third possible interval [^3] Used, for example, by the sacreBLEU package in machine translation evaluation. is to take half of the distance between the percentiles and report the interval estimate $\text{estimate}\pm\text{half-width}$.
The detailed results can be seen in Table 13 in the Appendix. The column F1 (true) indicates the observed F1 score, computed as an average over the $S$ seeds on the original data $\mathcal{D}$. All three types of CIs seem to consistently report similar ranges for the same estimator regardless of whether the languages are randomized. The bias and variance of the different estimators differ depending on the experimental condition. When the languages are kept constant, standard errors range from approximately 0.1 to 0.5 F1 points, except for XLM-R (ind.), where $\hat{\text{se}}(\hat{\mu}_{gm})=1.19$. When the languages are randomized, the standard errors increase considerably. All standard errors are now at least .5 F1 with many significantly larger. This is most likely related to increased variation in the normalization constant of the weights as the languages are randomized. Finetuning on all languages helps in most cases even with the languages randomized, both in terms of observed F1 scores and the observed variation. The exception to this is mBERT for which finetuning on all languages concatenated hurts performance.
@@
To obtain estimates of the pairwise differences in the overall population-level means, $\Delta_{ml}=\hat{\mu}^{(m)}_{am}-\hat{\mu}^{(n)}_{am}$, we compute a pairwise difference matrix for each replication $r$. Given the $L\times L\times R$-dimensional array of pairwise differences, we average over the $R$-dimension to obtain a single $L\times L$ table of average differences. This is displayed as the top table in Table 5. We also compute a standard deviation of each pairwise difference and divide the mean by the SD to obtain an estimate of the effect size. Looking at the second part of Table 5, all effect sizes except the differences between XLM-R (ind.) and Glot500 (ind.) as well as XLM-R (conc.) and XLM-R (ind.) appear significantly different from zero.
We also conduct another simulation where we subsample $10$ datasets without replacement from the original 61 and estimate effect sizes as above. The motivation for this simulation is to incorporate between-language variability into the estimation of overall performance. Intuitively, this also corresponds to a simulating the effect of downstream usage of the tested models on smaller sets of tasks. These results are shown in the bottommost part of Table 5. Overall, the effect size estimates are much smaller, with many shrinking below magnitude 2 due to the increased standard deviation. This suggests that using fixed benchmarks may understate the differences between two models.
@@
Table 5: Means and effect sizes of pairwise F1 score differences $\Delta_{ml}=\hat{\mu}^{(m)}_{am}-\hat{\mu}^{(n)}_{am}$ based on $R=5000$ simulations.
Ranks
@@
Table 6: Frequency distribution of ranks for each model based on $R=5,000$ bootstrap simulations of aggregate performance using the arithmetic mean $\hat{\mu}_{am}(\mathbf{x}^{(m)})$.
To estimate a distribution for model rankings, we rank each replication of the models based on the arithmetic mean across tasks and compute an empirical distribution approximation. The results of this simulation using are shown in Table 6. The bolded elements indicate the ranks observed using the original data. Overall, it seems like the ranks are stable for most models, except for Glot500 (ind.) and XLM-R (ind.) which compete for the 3rd place. We also repeated the simulation with fixed languages but observed that the ranks remained constant. This is in line with the observation from Section 6.1 that the between-language variation comprises by far the largest variance component. We suspect that under the “subsampling” condition, the ranks would display a lot more variability but leave this for future work.
@@ 7 Conclusion
In this paper, we have shown how replication and resampling can be helpful tools in the evaluation of multilingual NLP models. Through experiments on question answering, machine translation with LLMs, as well as NER using BERT/XLM-R-style PLMs, we showed how resampling-based inference can be useful in estimating variance components which can help researchers understand what their models are robust to and how much they might be underestimating total variation when computing standard errors. Since many hypothesis tests are based on computing standard errors, we urge researchers to estimate variance components to hedge against false positive discoveries.
We also showed how to approximate the distribution of score vectors using the parametric and nonparametric bootstrap, which allows us to estimate difficult-to-estimate parameters numerically, such as standard errors of pairwise differences or aggregate functions whose variance is unknown and rank distributions. We showed that in leaderboard-style aggregate evaluations, the between-task term may dominate the variance of the aggregate estimate.
Given these findings, we stress the importance of not relying on static leaderboards that, by construction, underestimate the total variability by neglecting the variance across tasks. We hope that this paper can stimulate further research into resampling-based evaluation in NLP that goes beyond hypothesis testing and computing error bars for individual performance scores.
@@ B.1 XQUAD
XQUAD (Artetxe et al., 2020) is a a multilingual extension of the SQuAD v1.1 question answering benchmark (Rajpurkar et al., 2016). XQUAD includes translations 1,190 questions, originally in English, into Arabic, German, Greek, Spanish, Hindi, Romanian, Russian, Thai, Turkish, Vietnamese, and Mandarin Chinese.
@@
Our analysis is based on the OpenNER 1.0 multilingual NER benchmark (Palen-Michel et al., 2024) which consists of 61 unique datasets covering a total of 51 languages. The authors experiment with three pretrained language models (PLM): mBERT (Devlin et al., 2019), XLM-R (Conneau et al., 2020), and Glot500-m (Imani et al., 2023). Each PLM is finetuned using two experimental conditions per language. In the “individual” condition, each PLM is finetuned separately on the train split of each of the 61 datasets. In the “concatenated” setting, the train splits of all 61 datasets are concatenated [^5] Due to computational considerations, before concatenation, the highest-resource datasets are downsampled. together before finetuning each of the three PLMs on it. The authors run each experiment 10 times using different seeds (values 42–51).
@@ C.1 Variance components
aya-expanse-8b TowerInstruct-Mistral-7B
Task TowerInstruct-Mistral-7B Clarus-7B Clarus-7B
@@
aya-expanse-8b 7.11 - - -
TowerInstruct-7B 8.49 - - -
Clarus-7B 4.50 - - -
@@
aya-expanse-8b 0.64 0.10 0.40 0.80
TowerInstruct-7B 0.77 0.22 0.28 1.15
Clarus-7B 0.70 0.23 0.22 1.12
@@
aya-expanse-8b 0.34 0.10 0.15 0.55
TowerInstruct-7B 0.47 0.16 0.13 0.78
Clarus-7B 0.48 0.20 0.14 0.92
@@
aya-expanse-8b 0.53 0.08 0.34 0.67
TowerInstruct-7B 0.60 0.17 0.25 0.93
Clarus-7B 0.49 0.13 0.17 0.70
@@
aya-expanse-8b 8.05 - - -
TowerInstruct-7B 8.99 - - -
Clarus-7B 9.66 - - -
@@
aya-expanse-8b 0.53 0.06 0.43 0.68
TowerInstruct-7B 0.48 0.08 0.33 0.64
Clarus-7B 0.67 0.16 0.34 0.99
@@
aya-expanse-8b 0.28 0.09 0.11 0.46
TowerInstruct-7B 0.25 0.08 0.08 0.42
Clarus-7B 0.43 0.18 0.11 0.82
@@
aya-expanse-8b 0.44 0.04 0.35 0.51
TowerInstruct-7B 0.40 0.04 0.32 0.48
Clarus-7B 0.50 0.06 0.32 0.58
@@
aya-expanse-8b 1.21 - - -
TowerInstruct-7B 2.44 - - -
Clarus-7B 5.75 - - -
@@
aya-expanse-8b 0.36 0.07 0.26 0.55
TowerInstruct-7B 0.38 0.10 0.23 0.63
Clarus-7B 0.74 0.12 0.53 1.01
@@
aya-expanse-8b 0.20 0.06 0.10 0.34
TowerInstruct-7B 0.22 0.11 0.08 0.52
Clarus-7B 0.50 0.15 0.27 0.87
@@
aya-expanse-8b 0.30 0.06 0.22 0.46
TowerInstruct-7B 0.30 0.06 0.21 0.44
Clarus-7B 0.53 0.06 0.41 0.66
@@
Glot500 (concat) 6.23 - - -
Glot500 (ind) 8.16 - - -
mBERT (concat) 14.28 - - -
mBERT (ind) 9.83 - - -
XLM-R (concat) 6.24 - - -
XLM-R (ind) 7.04 - - -
@@
Glot500 (concat) 1.76 0.93 0.51 4.24
Glot500 (ind) 1.49 0.52 0.68 2.74
mBERT (concat) 1.27 0.76 0.12 3.14
mBERT (ind) 1.26 0.75 0.53 3.53
XLM-R (concat) 1.58 1.24 0.48 5.45
XLM-R (ind) 1.47 1.11 0.37 4.80
@@
Glot500 (concat) 0.84 0.62 0.13 2.63
Glot500 (ind) 0.75 0.41 0.23 1.85
mBERT (concat) 0.71 0.57 0.04 2.34
mBERT (ind) 0.66 0.48 0.19 2.10
XLM-R (concat) 0.64 0.56 0.15 2.40
XLM-R (ind) 0.66 0.51 0.14 2.18
@@
Glot500 (concat) 1.54 0.81 0.47 3.72
Glot500 (ind) 1.28 0.40 0.59 2.18
mBERT (concat) 1.02 0.67 0.12 2.79
mBERT (ind) 1.05 0.60 0.46 2.87
XLM-R (concat) 1.13 1.09 0.40 4.77
XLM-R (ind) 1.30 0.74 0.33 3.28
@@
Clarus-7B TowerInstruct-7B aya-expanse-8b gemma-2-9b
@@
Arabic 0.37 0.66 0.76 0.57 0.60 0.83 0.27 1.08 1.11 0.91 0.92 1.29
German 0.35 0.64 0.73 0.83 0.86 1.20 0.51 1.13 1.24 0.60 1.10 1.25
Greek 0.34 0.84 0.91 0.42 0.51 0.66 1.00 1.02 1.43 0.82 1.01 1.30
English 0.65 1.13 1.30 0.31 0.98 1.02 0.86 1.01 1.33 0.85 1.17 1.44
Spanish 1.22 1.06 1.62 0.91 0.85 1.24 0.47 1.06 1.16 1.21 0.97 1.55
Hindi 0.14 0.83 0.84 0.20 0.37 0.42 0.59 1.13 1.28 0.63 1.03 1.21
Romanian 0.38 0.83 0.91 0.76 0.84 1.13 0.60 1.10 1.25 1.17 1.06 1.58
Russian 0.75 0.83 1.12 0.49 1.00 1.11 0.45 1.12 1.21 1.20 0.95 1.53
Thai 0.69 0.89 1.12 0.24 0.45 0.51 0.90 1.13 1.44 0.70 1.01 1.23
Turkish 0.18 0.57 0.60 0.27 0.60 0.66 0.77 1.11 1.35 0.63 1.07 1.24
Vietnamese 0.59 0.50 0.77 0.31 0.65 0.72 0.57 1.03 1.18 0.68 0.98 1.20
Chinese 0.57 0.70 0.90 0.88 0.80 1.19 0.56 1.27 1.39 0.56 1.02 1.16
@@
aya-expanse-8b TowerInstruct-7B Clarus-7B

### v2

@@ 4.2 Resampling for model comparison
We seek to quantify the uncertainty in the XQuAD F1 scores in each language as well as aggregated across languages. Specifically, we compute pairwise model differences on each language as well as between aggregate scores computed using the arithmetic mean, geometric mean and median. We also compute rankings as point estimates.
We estimate the difference between models using the following parametric bootstrap resampling procedure. Given a language, we first take an estimate of each model’s average performance on it by averaging the $S=5$ observed values. We then use the estimate of the within-language SD, $\eta_{ml}$, to get a randomly resampled performance score for each model.
We estimate $\eta_{ml}$ by first computing the standard deviation of scores across seeds and then use the average bootstrap SD $\tau_{ml}$ across the $S=5$ values computed by lm-evaluation-harness to form an aggregate SD estimate of $\eta_{ml}$. We then sample a noise term using a random draw from $\varepsilon\sim\mathcal{N}(0,\eta_{ml}^{2})$ yielding a resampled score $x+\varepsilon$ with variance $\eta_{ml}^{2}=\sigma_{ml}^{2}+\tau_{ml}^{2}$. While the score itself may not be Gaussian, we believe that modeling the noise terms as such is reasonable, as we observe mostly small score deviations across both seeds and bootstrap resamples.
We repeat this process $R=1000$ times. For each replication, we compute a performance difference on each language, and an aggregated score of performance scores across languages, using the arithmetic mean, geometric mean and median. We also rank the models based on the randomly replicated performance scores. The results of this procedure are summarized in several tables below.
Table 2 summarizes pairwise differences between models and their associated SDs. Table 3 gives the rank distribution of each model using each of the summary statistics. While Aya Expanse 8B and TowerInstruct 7B reliably place first and last regardless of metric, the rank distributions of Clarus 7B and Gemma 2 9B can vary significantly depending on which aggregation function is used.
@@
5 Task 2: Machine translation with LLMs
@@ 5.1 Variance components
The lm-evaluation-harness library we use to run our experiments automatically computes bootstrapped standard errors for BLEU, ChrF++ and COMET. In addition to the bootstrap SE, we incorporate model-side uncertainty by sampling 5 translation hypotheses for each source sentence from each model. The summarized results are shown in Table 8 in the Appendix. A detailed variance decomposition of the within-language variation is shown in Table 11 in the Appendix. As with XQuAD, the between-language variability ($\nu$) is much larger than either the model-side or bootstrap uncertainty. Most variance components are also less than 1.0 BLEU or ChrF++ points.
@@ 5.2 Resampling for model comparison
As with XQuAD, we also wish to quantify the uncertainty in the performance differences between the two models on each language. Given observed BLEU/ChrF++/COMET scores for each language, we can easily compute a point estimate by subtraction. We quantify the uncertainty in this point estimate by resampling as with XQuAD. The estimated differences are displayed in Table 7. The most obvious observation is that Clarus 7B clearly underperforms relative to Aya Expanse 8B and TowerInstruct 7B. The latter two models perform more similarly to each other and significant differences were found in 6 out of 16 translation using BLEU. Out of the 6 significant differences, TowerInstruct 7B beat Aya Expanse 8B 4 out of 6 times. When using ChrF++, significant differences were found in 10 out of 16 translation tasks, with TowerInstruct 7B leading all of them.
@@ 6 Task 3: NER
As a third task, we show how resampling can be used for model comparison in multilingual NER when finetuning pretrained language models. Our analysis is based on the OpenNER 1.0 multilingual NER benchmark (Palen-Michel et al., 2025) which consists of 61 unique datasets covering a total of 51 languages. Since we reanalyze the results of the original paper, we use the same models: mBERT and XLM-R. For each model architecture, we vary the training data according to two conditions on each language-dataset pair: (i) individual, where each model is finetuned with data from only a given training set and (ii) concatenated, where each model is finetuned with a concatenated and downsampled version of all the training data included in OpenNER. In addition to the original results, we construct $B=100$ bootstrap replicated data sets and reuse the same predictions from the original test set. This gives us a total of $R=SB=1000$ replicated scores for each of the 61 datasets.
@@ 6.1 Variance components via resampling
Like with QA and MT, we investigate how resampling-based inference can be used to better make sense of experimental variation by estimating the variance components due to model and data-side uncertainty for each of the 61 languages pairs that comprise the overall benchmark. A concise display of inferences for $\nu$, $\sigma$ and $\tau$ are shown in Table 4. A more detailed table of estimates is shown in the Appendix in Table 12.
@@
Glot500 1.59(2.51) 0.61(0.39)
$\sigma_{l}$ mBERT 1.15(2.55) 0.81(0.45)
XLM-R 1.24(3.53) 0.64(0.42)
Glot500 1.27(0.95) 1.17(0.79)
$\tau_{l}$ mBERT 1.36(0.86) 1.47(0.91)
XLM-R 1.23(0.84) 1.18(0.79)
Table 4: Summary of variance components for NER. The three groups of rows, $\nu$, $\sigma$ and $\tau$ correspond to between-language, seed, and bootstrap SDs. Subscripts indicate standard deviations across languages.
Looking at Table 4, we see that both model and data-side variability tend to lie in the 0.6–1.6 F1 range. When we compare models in the individual versus the concatenated conditions, the models in the concatenated condition exhibit distinctly lower model-side variability. This suggests that languages with less training data benefit from larger, multilingual finetuning, even if no positive transfer learning is taking place, possibly due to a reduction in gradient instability.
Model-side variability, on the other hand, remains relatively constant across individual and concatenated conditions which suggests that the size of finetuning sets may not help much in terms of dataset-to-dataset generalization. This observation also highlights how the model and data-side variance components measure fundamentally different aspects of performance variability and how both are needed for thorough performance assessment.
@@ 6.2 Resampling for model comparison
As with QA and NMT, we also estimate quantities other than variance components. To estimate uncertainties for model rankings and pairwise differences between models, we resample the $R=SB=1000$ replications within each language. Unlike our QA and MT experiments, we do this entirely nonparametrically by sampling a seed and replication index at random. We also show how the between-language score variability ($\nu$) can be incorporated into the analysis by also resampling what languages are considered when estimating overall performance.
@@
Table 5: Means and effect sizes of pairwise F1 score differences $\Delta_{m}=\hat{\mu}^{(m)}_{\text{arith}}-\hat{\mu}^{(n)}_{\text{arith}}$.
To obtain estimates of the pairwise differences of average performance between models, $\Delta_{mn}=\hat{\mu}^{(m)}_{\text{arith}}-\hat{\mu}^{(n)}_{\text{arith}}$, we compute a pairwise difference matrix for each replication $r$. Given the $M\times M\times R$-dimensional array of pairwise differences, we average over the $R$-dimension to obtain a single $M\times M$ table of average differences. This is displayed as the top table in Table 5.
To account for uncertainty, we also compute a standard deviation of each pairwise difference across the $R$-dimension. We then divide the estimated mean difference by the SD to obtain an estimate of the effect size which allows us to assess the statistical significance of the observed differences. These effect sizes are displayed in the second part of Table 5. Except for the differences between XLM-R (ind.) and Glot500 (ind.), as well as XLM-R (conc.) and XLM-R (ind.), all estimates appear to be significantly different from zero.
We also conduct another simulation where we subsample $10$ datasets without replacement and otherwise estimate effect sizes as above. The motivation for this is to incorporate between-language variability into the estimation of overall performance. Intuitively, this also corresponds to a simulating the effect of downstream usage of the tested models on smaller sets of tasks. These results are shown in the bottom part of Table 5. Overall, the effect size estimates are much smaller, with many shrinking below magnitude 2 due to the increased standard error. This suggests that using fixed benchmarks may understate the differences between two models . While the chosen threshold for significance is arbitrary, the pattern of effect size shrinkage due to the increased variability is robust.
@@
Table 6: Bootstrap-resampled frequency distribution of ranks for each model using arithmetic mean as an aggregation function.
Ranks
To estimate a distribution for model rankings, we recompute the model ranks using the arithmetic mean across languages for each replication. This yields an empirical distribution estimate of model ranks which is displayed in Table 6. The bolded elements indicate the ranks observed using the original data. Overall, it seems like the ranks are stable for most models, except for Glot500 (ind.) and XLM-R (ind.) which compete for the 3rd place. We also repeated the simulation with fixed languages but observed that the ranks remained constant. This is in line with the observation from Section 6.1 that the between-language variation comprises by far the largest variance component. We suspect that under the “subsampling” condition, the ranks would display a lot more variability but leave this for future work.
@@ 7 Conclusion
In this paper, we have shown how replication and resampling can be helpful tools in the evaluation of multilingual NLP models. Through experiments on question answering, machine translation, and named entity recognition, we showed how resampling-based inference can be used to estimate different variance components which describe how robust model performance may be to model-related randomness as well as test set composition and between-language variation. We also showed that the between-language component may dominate the standard error of estimators such as the arithmetic mean in leaderboard-style evaluation settings that involve multiple languages/tasks.
In addition to variance components, we also showed how resampling can be used to estimate the distributions and standard errors of quantities with no closed form expressions, such as pairwise differences between models and rank distributions. We also explained how standard errors relate to statistical significance and showed that underestimated SEs may lead to overly optimistic results that ultimately fail to replicate.
We also wish to stress stress that the techniques introduced in this paper need not translate to significant additional computational overhead. In particular, between-language/task variance can be computed without any resampling and within-test set bootstrap resampling procedures can be run in seconds on a modern consumer CPU. While model-side variance may require comparatively more resources in some cases, training on multiple seeds is often only relevant when working with smaller, non-LLM models. With LLMs, it is sufficient to sample multiple responses for a given input which can be done at the fraction of the cost of an entire retraining run.
It is our hope that this paper can stimulate further research into understanding the sources, extent, and nature of performance variation in empirical NLP research. We conclude with a vision for uncertainty-aware multilingual/multitask evaluation in NLP research. Researchers should:
  1. 1.
    Thoughtfully consider all relevant sources of randomness when evaluating and comparing different models against each other. Draw more than one set of predictions to understand how variable the performance of a model is, and use bootstrapping to complement model-side uncertainty with data-side uncertainty estimates.
  2. 2.
    In multilingual/multitask evaluation setups, strive to understand how observed differences between models vary from task to task. Incorporate estimates of between-language variation into statistical comparisons to avoid false positives that may not be replicable.
  3. 3.
    If model performance or observed differences are very noisy, communicate this clearly and describe the differences between tasks. This is preferable to attempting to distill results to a binary judgment about which model is better. If feasible, show example cases where models agree/disagree.
  4. 4.
    Use resampling with the replications obtained from the above steps to estimate distributions and associated uncertainty for downstream quantities of interest, such as pairwise differences and model rankings.
@@ B.1 XQuAD
XQuAD (Artetxe et al., 2020) is a multilingual extension of the SQuAD v1.1 question answering benchmark (Rajpurkar et al., 2016). XQuAD includes translations 1,190 questions, originally in English, into Arabic, German, Greek, Spanish, Hindi, Romanian, Russian, Thai, Turkish, Vietnamese, and Mandarin Chinese.
@@
Our analysis is based on the OpenNER 1.0 multilingual NER benchmark (Palen-Michel et al., 2025) which consists of 61 unique datasets covering a total of 51 languages. The authors experiment with three pretrained language models (PLM): mBERT (Devlin et al., 2019), XLM-R (Conneau et al., 2020), and Glot500-m (Imani et al., 2023). Each PLM is finetuned using two experimental conditions per language. In the “individual” condition, each PLM is finetuned separately on the train split of each of the 61 datasets. In the “concatenated” setting, the train splits of all 61 datasets are concatenated [^3] Due to computational considerations, before concatenation, the highest-resource datasets are downsampled. together before finetuning each of the three PLMs on it. The authors run each experiment 10 times using different seeds (values 42–51).
@@ C.1 Variance components
Aya Expanse 8B TowerInstruct 7B
Task TowerInstruct 7B Clarus 7B Clarus 7B
@@
Aya Expanse 8B 7.11 - - -
TowerInstruct 7B 8.49 - - -
Clarus 7B 4.50 - - -
@@
Aya Expanse 8B 0.64 0.10 0.40 0.80
TowerInstruct 7B 0.77 0.22 0.28 1.15
Clarus 7B 0.70 0.23 0.22 1.12
@@
Aya Expanse 8B 0.34 0.10 0.15 0.55
TowerInstruct 7B 0.47 0.16 0.13 0.78
Clarus 7B 0.48 0.20 0.14 0.92
@@
Aya Expanse 8B 0.53 0.08 0.34 0.67
TowerInstruct 7B 0.60 0.17 0.25 0.93
Clarus 7B 0.49 0.13 0.17 0.70
@@
Aya Expanse 8B 8.05 - - -
TowerInstruct 7B 8.99 - - -
Clarus 7B 9.66 - - -
@@
Aya Expanse 8B 0.53 0.06 0.43 0.68
TowerInstruct 7B 0.48 0.08 0.33 0.64
Clarus 7B 0.67 0.16 0.34 0.99
@@
Aya Expanse 8B 0.28 0.09 0.11 0.46
TowerInstruct 7B 0.25 0.08 0.08 0.42
Clarus 7B 0.43 0.18 0.11 0.82
@@
Aya Expanse 8B 0.44 0.04 0.35 0.51
TowerInstruct 7B 0.40 0.04 0.32 0.48
Clarus 7B 0.50 0.06 0.32 0.58
@@
Aya Expanse 8B 1.21 - - -
TowerInstruct 7B 2.44 - - -
Clarus 7B 5.75 - - -
@@
Aya Expanse 8B 0.36 0.07 0.26 0.55
TowerInstruct 7B 0.38 0.10 0.23 0.63
Clarus 7B 0.74 0.12 0.53 1.01
@@
Aya Expanse 8B 0.20 0.06 0.10 0.34
TowerInstruct 7B 0.22 0.11 0.08 0.52
Clarus 7B 0.50 0.15 0.27 0.87
@@
Aya Expanse 8B 0.30 0.06 0.22 0.46
TowerInstruct 7B 0.30 0.06 0.21 0.44
Clarus 7B 0.53 0.06 0.41 0.66
@@
Glot500 (concat) 6.30 - - -
Glot500 (ind) 8.27 - - -
mBERT (concat) 14.38 - - -
mBERT (ind) 9.95 - - -
XLM-R (concat) 6.31 - - -
XLM-R (ind) 7.13 - - -
@@
Glot500 (concat) 1.33 0.86 0.33 4.88
Glot500 (ind) 2.19 2.56 0.39 13.98
mBERT (concat) 1.70 0.99 0.24 5.42
mBERT (ind) 1.90 2.61 0.14 20.62
XLM-R (concat) 1.36 0.87 0.15 5.08
XLM-R (ind) 1.93 3.54 0.13 28.12
@@
Glot500 (concat) 0.61 0.39 0.19 2.09
Glot500 (ind) 1.59 2.51 0.16 13.69
mBERT (concat) 0.81 0.45 0.12 2.29
mBERT (ind) 1.15 2.55 0.06 20.35
XLM-R (concat) 0.64 0.42 0.08 2.46
XLM-R (ind) 1.24 3.53 0.04 28.12
@@
Glot500 (concat) 1.17 0.79 0.13 4.79
Glot500 (ind) 1.27 0.95 0.12 5.31
mBERT (concat) 1.47 0.91 0.21 5.33
mBERT (ind) 1.36 0.86 0.13 5.20
XLM-R (concat) 1.18 0.79 0.13 5.00
XLM-R (ind) 1.23 0.84 0.12 5.22
@@
Clarus 7B TowerInstruct 7B Aya Expanse 8B Gemma 2 9B
@@
Arabic 0.20 0.52 0.56 0.46 0.43 0.63 0.25 1.15 1.18 0.48 0.80 0.93
Chinese 0.36 0.55 0.66 0.52 0.63 0.82 0.65 1.28 1.43 0.51 0.98 1.10
English 0.93 1.27 1.58 0.60 1.05 1.21 0.96 1.41 1.71 0.95 1.18 1.51
German 0.57 0.54 0.78 0.74 0.81 1.10 0.79 1.31 1.53 0.69 1.05 1.26
Greek 0.36 0.74 0.82 0.17 0.33 0.37 0.71 1.07 1.29 0.97 0.89 1.31
Hindi 0.32 0.77 0.83 0.12 0.15 0.19 0.89 1.34 1.61 0.72 0.96 1.20
Romanian 0.36 0.80 0.88 0.51 0.76 0.91 0.34 1.33 1.37 0.85 1.01 1.32
Russian 0.88 0.79 1.18 0.31 0.95 1.00 0.53 1.21 1.32 0.80 0.84 1.16
Spanish 1.25 1.11 1.67 0.49 0.76 0.91 0.88 1.26 1.54 1.09 0.88 1.40
Thai 0.59 0.70 0.91 0.19 0.30 0.35 0.72 1.06 1.28 0.65 0.92 1.13
Turkish 0.16 0.37 0.40 0.40 0.41 0.57 1.23 1.21 1.73 0.67 1.00 1.20
Vietnamese 0.34 0.24 0.42 0.42 0.52 0.67 1.18 1.19 1.67 0.51 0.89 1.03
@@
Aya Expanse 8B TowerInstruct 7B Clarus 7B